# Artificial Hippocampus Networks for Efficient Long-Context Modeling

### v1

@@ 4.5 Probing AHN with gradients visualization
where $f^{\prime}(\cdot)$ and $f(\cdot)$ denote the teacher and student forward models, $h_{\text{AHN}}$ represents the compressed memory of AHN, $x_{\text{win}}$ are in-window token embeddings, and $x_{\text{out}}$ are out-of-window embeddings. Out-of-window tokens with small gradient magnitudes indicate that their information has already been well captured in AHN’s compressed memory. As illustrated by the math example in Figure 5, AHN tends to preserve the information of mathematical symbols and numbers while neglecting less critical ones such as pronouns and special tokens, demonstrating its effectiveness as a targeted compression module.
@@ 5 Conclusion and discussion
Limitations and future works. While AHNs strike an effective balance between computational efficiency and memory fidelity, their fixed-size compressed memory inevitably entails some information loss and may impair performance on tasks that require exact recall, as detailed in the appendix. Furthermore, since our study adopts a parameter-efficient self-distillation setup, performance remains capped by the underlying base models’ capacity. Future work may explore stronger recall mechanisms and full-parameter training to further unlock the potential of AHNs. For application scenarios, the AHN framework opens up opportunities in long-context domains with sparse information or constrained resources, such as lifelong learning, streaming video processing, and deployment on edge devices.
@@ Acknowledgement
We thank Shi Guang, Haoqi Fan, Tianle Cai, Deyao Zhu, Tenglong Ao, Ge Zhang, Wenhao Huang, and Liang Xiang for valuable discussions.

### v2

@@ 3.5 Probing AHN with gradients visualization
where $f^{\prime}(\cdot)$ and $f(\cdot)$ denote the teacher and student forward models, $h_{\text{AHN}}$ represents the compressed memory of AHN, $x_{\text{win}}$ are in-window token embeddings, and $x_{\text{out}}$ are out-of-window embeddings. Out-of-window tokens with small gradient magnitudes indicate that their information has already been well captured in AHN’s compressed memory.
Figure 5 shows an example from AceMath-Instruct-Training-Data [57]. The full sequence has 811 tokens, and we evaluate the AHN-augmented model using a 512-token sliding window (AHN activates once the context surpasses 512 tokens). The snippet shows the gradients for the first 139 tokens. As illustrated in Figure 5, AHN tends to preserve the information of mathematical symbols and numbers while neglecting less critical ones such as pronouns and special tokens, demonstrating its AHNs can learn to prioritize more informative tokens for storage.
4 Related work
4.1 Memory in neural networks
Memory mechanisms play a crucial role in enabling neural networks to process and retain information over time, which is essential for tasks that require understanding of temporal dependencies, sequential data, or context preservation. Traditional feedforward neural networks lack the capability to maintain information across time steps, which limits their effectiveness in tasks such as language modeling, sequence prediction, and reasoning. To address this limitation, Recurrent Neural Networks (RNNs) are introduced [92, 43, 25, 37, 38]. RNNs maintain a hidden state that is updated at each time step, allowing information to persist across sequences. However, vanilla RNNs suffer from issues such as vanishing and exploding gradients, making it difficult to capture long-term dependencies [11]. To mitigate these problems, more advanced architectures like Long Short-Term Memory (LSTM) networks [35] and Gated Recurrent Unit (GRU) [15] are proposed. These models incorporate gating mechanisms that regulate the flow of information, enabling them to learn longer-term dependencies more effectively. Because these RNN-like models maintain a fixed-size memory and a consistent memory update cost for each input token, they are highly efficient for processing long sequences. Therefore, our AHNs are designed within the RNN paradigm to inherit this advantageous property.
4.2 Memory management
RNN-like models [25, 35, 15, 45, 69, 81, 31, 103, 115, 104, 19, 9, 105] maintain memory through a fixed-size hidden state, regardless of input sequence length. Therefore, memory caching is not a major concern for these architectures. In contrast, Transformers store key-value (KV) pairs for every token in the input sequence, resulting in linear growth of the KV cache with sequence length. This results in significant memory consumption and presents a major challenge for processing long sequences. To mitigate this issue, various approaches have been proposed [49], including KV cache selection [28, 51, 98, 33, 116, 56, 1, 95, 84], budget allocation [13, 101, 26, 97], merging [63, 86, 89, 55], quantization [106, 79, 36, 96, 53, 77], low-rank decomposition [108, 22], external memory [67, 88], and neural architecture design [78, 2, 54, 40, 82, 107, 94, 62]. Among them, a straightforward strategy is to use a sliding window for attention [85], but this method discards KV pairs outside the window, thereby losing long-range context. Sparse Transformers [14] address this by retaining KV pairs at specific pattern positions to capture long-range dependencies, but still drop portions of the KV cache, potentially missing important information. Transformer-XL [18] introduces a segment-level recurrence mechanism by caching the last segment of hidden states as a First-In, First-Out (FIFO) memory. Compressive Transformer [73] extends this by compressing older memories into a secondary FIFO memory, but it still discards memory once the slots are full. In contrast, AHNs adopt an RNN-like paradigm that continually compresses KV pairs outside the sliding window into a lifelong compressed memory, rather than discarding them outright [52, 62, 74]. AHNs (like AHN-GDN [105]) can also dynamically control memory decay [19, 75, 104, 105]. Recent studies integrate RNNs and attention either in interleaved layers [52, 74, 19, 105, 48] or within a single layer [62, 10, 50]. By contrast, we abstract the compression module as an AHN concept, yielding a more general memory framework. We employ a sliding-window attention mechanism, activating AHNs whenever a token leaves the window. Additionally, we introduce a simple self-distillation scheme that trains AHNs efficiently.
Compared with recent attention-RNN-Hybrid works [62, 87, 112] and concurrent work [41], the goal of the AHN memory framework is to leverage the efficiency of RNNs specifically to address the computational bottleneck of attention on extra-long sequences. The distinct insight introduced by AHNs is to employ a large sliding-window size (e.g., 32k) for attention, such that the RNN-like AHN modules only activate when the sequence length exceeds this window. This design provides two major advantages: 1) It highlights the efficiency benefits of RNNs on extra-long context tasks (e.g., 128k), achieving substantial FLOP and memory-cache savings. In contrast, for short-context tasks where attention is already efficient, introducing RNNs provides no efficiency gain. 2) It requires no additional effort to preserve attention performance on short-context tasks, because AHNs remain inactive in these regimes, and the model behaves exactly as a pure attention-based Transformer.
@@ 5 Conclusion and discussion
Limitations and future works. While AHNs strike an effective balance between computational efficiency and memory fidelity, their fixed-size compressed memory inevitably entails some information loss and may impair performance on tasks that require exact recall, as detailed in the appendix. Future work may explore stronger recall mechanisms. For application scenarios, the AHN framework opens up opportunities in long-context domains with sparse information or constrained resources, such as lifelong learning, streaming video processing, and deployment on edge devices.
@@ Acknowledgement
We thank Shi Guang, Haoqi Fan, Tianle Cai, Deyao Zhu, Tenglong Ao, Shilong Liu, Ge Zhang, Wenhao Huang, and Liang Xiang for valuable discussions.